1/2

Heinrich Heine,  Heinrich Heine,
The Lorelei  The Lorelei
tradukita de Mark Twain  tradukita de L. W. Garnham

An ancient legend of the Rhine

I cannot divine what it meaneth, This haunting nameless pain: A tale of the bygone ages Keeps brooding through my brain:

I do not know what it signifies.

That I am so sorrowful?

A fable of old Times so terrifies, Leaves my heart so thoughtful.

The faint air cools in the gloaming, And peaceful flows the Rhine, The thirsty summits are drinking The sunset’s flooding wine; The air is cool and it darkens, And calmly flows the Rhine; The summit of the mountain hearkens In evening sunshine line.

The loveliest maiden is sitting High-throned in yon blue air, Her golden jewels are shining, She combs her golden hair; The most beautiful Maiden entrances Above wonderfully there, Her beautiful golden attire glances, She combs her golden hair.

She combs with comb that is golden, And sings a weird refrain That steeps in a deadly enchantment The listener’s ravished brain:

With golden comb so lustrous, And thereby a song sings, It has a tone so wondrous, That powerful melody rings.

The doomed in his drifting shallop, Is tranced with the sad sweet tone, He sees not the yawing breakers, He sees but the maid alone:

The shipper in the little ship It effects with woe sad might; He does not see the rocky slip, He only regards dreaded height.

…  …

2/2

The pitiless billwos engulf him!- So perish sailor and bark; And this, with her baleful singing, Is the Lorelei’s gruesome work.

I believe the turbulent waves

Swallow the last shipper and boat; She with her singing craves All to visit her magic moat.

Traduko de la Germana poemo “Die Lorelei” de Heinrich Heine (∗1797-12-13 – †1856-02-17) en la Anglan de Mark Twain.

Arg-2-11 (2003-10-13 04:42:59)

A Tramp Abroad. Vol 1-2. Leibzig: Tauchnitz, 1880 Band I, Mark Twain 1880 (rf. http: // www. loreley. com/ loreley/ marctwai. htm )

Traduko de la Germana poemo “Die Lorelei” de Heinrich Heine (∗1797-12-13 – †1856-02-17) en la Anglan de L. W. Garnham.

Arg-2-14 (2003-10-11 23:04:57)

L.W. Garnham, Bachelor of Arts, LEGENDS OF THE RHINE; mentioned in: A Tramp Abroad. Vol 1-2. Leibzig: Tauchnitz, 1880 Band I, Mark Twain 1880

http://www.poezio.net/version?poem-id=2&version-id=,11,14